# The Nonlinear Filter Model of Stream Cipher Redivivus

### v5

@@ 1 Introduction
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to map $(\mathbf{X},\mathbf{Y})$ to $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of both $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $h$. Assume that $\pi$ is chosen to be the bit reversal permutation. It is known [24] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, sufficient (fast) algebraic immunity when the number of variables is large enough (to protect against fast algebraic attacks at a specified security level), and at the same time is quite efficient to implement. Indeed, the (fast) algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers.
@@
We propose to use filtering function on a large number of variables. On the other hand, due to efficiency considerations, it is not possible to make the length of the LFSR too long. In particular, in all our concrete proposals, the ratio $n/L$ is close to half and is much higher than what was used in all previous proposals. Using a high value of $n/L$ means that the tap positions are placed much closer together than in previous proposals. This creates possibility of an overlap in the two sets of state bits determining two nearby keystream bits. Such overlap has the potential to cancel out terms and reduce the overall “complexity”. Our choice of tap positions is motivated by the requirement of ensuring that such cancellations do not take place. The generation of a keystream bit requires a call to the majority function. We set one of the goals of choosing the tap positions to be to ensure that the inputs to the two calls to majority for generating two different keystream bits have only a small overlap. In particular, we prove that for our choice of tap positions, there is no cancellation of terms of the majority function corresponding to two keystream bits. We evolved this design criterion for the tap positions in response to an attack [6] on an earlier version, and based on informal attack ideas (though not actual attacks) suggested by Subhadeep Banik, Willi Meier, and Bin Zhang also on a previous version. Further, we choose all the tap positions for the variables in $\mathbf{X}$ to the left of all the tap positions for the variables in $\mathbf{Y}$. This choice combined with the choice of $\pi$ as the bit reversal permutation allows us to prove that the quadratic terms arising from the application of the filtering function to the state bits corresponding to two different states do not cancel out with each other.
@@ 2 Preliminaries
The fast algebraic attack (FAA) was introduced in [22]. The idea of the attack is based on the following observation. Let $f$ be an $n$-variable function and suppose $g$ is another $n$-variable function of degree $e$ such that $gf$ has degree $d$. If both $e$ and $d$ are small, then $f$ is susceptible to an FAA. Given $f$, and for $e$ and $d$ satisfying $e+d\geq n$, it is known [22] that there exists functions $g$ and $h$ with ${\sf deg}(g)=e$ and ${\sf deg}(h)\leq d$ such that $gf=h$. Based on this result, we provide the following definition. For each $e\in\{1,\ldots,{\sf AI}(f)-1\}$, let $d\leq n-1-e$ be the maximum integer such that there do not exist $n$-variable functions $g$ and $h$ with ${\sf deg}(g)=e$, ${\sf deg}(h)=d$ and $gf=h$. We call the list of all such pairs $(e,d)$ to be the FAA-profile of $f$.
The fast algebraic immunity (FAI) of $f$ is a combined measure of resistance to both algebraic and fast algebraic attacks: ${\sf FAI}(f)\allowbreak=\allowbreak\min\left(2{\sf AI}(f),\allowbreak\min_{g\neq 0}\allowbreak\{{\sf deg}(g)\allowbreak+\allowbreak{\sf deg}(fg):\allowbreak 1\allowbreak\leq\allowbreak{\sf deg}(g)\allowbreak<\allowbreak{\sf AI}(f)\}\right)$.
So ${\sf FAI}(f)\allowbreak=\allowbreak\min(\allowbreak 2{\sf AI}(f),\allowbreak\min\{e+d+1\}\allowbreak)$, where the second minimum is taken over all pairs $(e,d)$ in the FAA-profile of $f$. Further, for any function $f$, $1+{\sf AI}(f)\leq{\sf FAI}(f)\leq 2\,{\sf AI}(f)$.
@@ 3 Construction from Maiorana-McFarland Bent Functions
Suppose that $\pi$ is a bit permutation, i.e. $\pi(X_{1},\ldots,X_{m})=(X_{\rho(1)},\ldots,X_{\rho(m)})$, for some permutation $\rho$ of $\{1,\ldots,m\}$. Then
@@
Suppose now that $\pi$ is a bit permutation. Then $\ell(\mathbf{X})=\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle$ is a linear function. If $w=0$, then $\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})=h(\mathbf{X})$, and so we have the result. So suppose $w>0$. From Lemma 1 of [16], we have ${\sf AI}(\ell(\mathbf{X})\oplus h(\mathbf{X}))\geq{\sf AI}(h(\mathbf{X}))-1$. So
@@
From Theorem 2, choosing $\pi$ to be any bit permutation results in the AI of ${\sf MM}_{2m}$ to be lower bounded by the AI of $h$. We make the following concrete choices.
@@ 4.2 Security Analysis
The sequence $w_{0},w_{-1},w_{-2},\ldots$ is known. If we guess the values of $s_{L-1},\ldots,s_{L-2\kappa+1}$ (a total of $\chi$ bits), then we obtain $s_{L-2\kappa}$ from the first equation, using the obtained value of $s_{L-2\kappa}$ in the second equation, we obtain the value of $s_{L-2\kappa-1}$, using the obtained value of $s_{L-2\kappa-1}$ in the third equation, we obtain the value of $s_{L-2\kappa-2}$, and so on. In other words, by guessing the values of $\chi$ of the state bits, we can obtain values of $L-\chi$ previous state bits, giving us the values of all $L$ bits of the state at a point $L-\chi$ steps in the backward direction. Once the complete state is known, the original guess can be verified by generating the keystream in the forward direction and matching with the obtained keystream. The method requires $2^{\chi}$ guesses, and for each guess about $Lm$ operations are required. By our choice of the tap positions for $\mathbf{X}$ and $\mathbf{Y}$, the value of $\chi$ is $12\kappa-$ 1. So the number of guesses required to successfully mount the attack is far greater than $2^{\kappa}$. This proves the resistance of the stream cipher to the inversion attack. Note that the inversion attack is prevented due to $\chi$ being sufficiently large. This was one of the countermeasures to inversion attack that was already proposed in [27].
@@
Let as above the state of the LFSR at some point be $\mathbf{s}^{(0)}$ and the subsequent states of the LFSR be denoted as $\mathbf{s}^{(1)},\mathbf{s}^{(2)},\ldots,\mathbf{s}^{(c-1)}$. For $i=0,\ldots,c-1$, let $\mathbf{t}^{(i)}={\sf proj}(\mathbf{s}^{(i)})$ and $w_{i}=f_{2m+1}(\mathbf{t}^{(i)})$. Recall from the definition of $\delta$ that for $10\leq i<j\leq c-$ 1, $\mathbf{t}^{(i)}$ and $\mathbf{t}^{(j)}$ have an overlap of at most $\delta$ bits. So in the $c$ $n$-bit strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ there are at least $nc-\delta c(c-1)/2$ unknown bits and hence there are at least $2^{nc-\delta c(c-1)/2}$ possibilities for these $c$ strings. Since $f_{2m+1}$ is balanced, knowledge of $w_{0}$ reduces the number of possibilities for $\mathbf{t}^{(0)}$ by a factor of half, and more generally, the knowledge of $w_{0},\ldots,w_{c-1}$ reduces the number of possibilities for $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ by a factor of $2^{c}$. So at least $2^{(n-1)c-\delta c(c-1)/2}$ possibilities remain for the $c$ strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$. For each of these possibilities, we obtain a system of $L$ equations in $L$ variables, the solution of which provides the value of initial state of the LFSR. The correctness of the obtained value of the state can be determined by generating the keystream from the obtained value and matching with the actual keystream. Solving the system of $L$ linear equations requires $L^{3}$ arithmetic operations, and generating the keystream from the obtained state require about $Ln$ operations. So the complexity of the attack is $2^{(n-1)c-\delta c(c-1)/2}(L^{3}+Ln)$ operations. Suppose the following condition holds.
@@
For a filtering function $f$ having ${\sf AI}(f)=a$, an algebraic attack requires about $N=\sum_{i=0}^{a}{L\choose i}$ keystream bits and has time complexity $T$ to be about $N^{\omega}$, where $\omega$ is the exponent of matrix multiplication (see Section 3.1.5 of [14]). We assume $\omega=2.8$. From Proposition 4, we assume that $a={\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})=\lceil m/2\rceil$, i.e. the algebraic immunity is equal to its lower bound. Let
@@
For any $(e,d)$ in the FAA-profile of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m}$, the Berlekamp-Massey step in a fast algebraic attack on the filter generator model has time complexity $\mathcal{O}(ED\log D)$, where $E=\sum_{i=0}^{e}{L\choose i}$ and $D=\sum_{i=0}^{d}{L\choose i}$ (see [31] and Section 3.1.5 of [14]). This complexity dominates the overall time complexity of a fast algebraic attack. Ignoring the logarithm term, we take $T$ to be equal to $ED$. The maximum value of $e$ is one less than the algebraic immunity $a$. The number $N$ of keystream bits required is about $2E$. From Proposition 4, we again assume that $a={\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})=\lceil m/2\rceil$. Define
$\gamma$ $=$ $\min_{\stackrel{{\scriptstyle 1\leq e\leq a-1,}}{{d=a-e}}}\left(\sum_{i=0}^{e}{L\choose i}\right)\left(\sum_{i=0}^{d}{L\choose i}\right).$ (25)
@@ 5 Efficiency of Computing ${\sf MM}_{n}$
Using the above algorithm, it is easy to show that the computation of the weight of an $m$-bit string, when $m=2^{r}-1$ with $r\geq 1$, requires $12^{r}-r-$ 1 full adders. From this it easily follows that the number of full adders required to compute the weight of an $m$-bit string for arbitrary $m$ is $O(m)$. Since a full adder can be implemented using a constant number of NAND gates, the number of NAND gates required to compute the weight of an $m$-bit string is also $O(m)$.

### v6

@@ 1 Introduction
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to map $(\mathbf{X},\mathbf{Y})$ to $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of both $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $h$. Assume that $\pi$ is chosen to be an affine map, i.e. the coordinate functions of $\pi$ are affine functions of its input variables. It is known [24] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, sufficient (fast) algebraic immunity when the number of variables is large enough (to protect against fast algebraic attacks at a specified security level), and at the same time is quite efficient to implement. Indeed, the (fast) algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers. To the best of our knowledge, our result is the first construction of bent functions with a provable lower bound on algebraic immunity.
@@
We propose to use filtering function on a large number of variables. On the other hand, due to efficiency considerations, it is not possible to make the length of the LFSR too long. In particular, in all our concrete proposals, the ratio $n/L$ is close to half and is much higher than what was used in all previous proposals. Using a high value of $n/L$ means that the tap positions are placed much closer together than in previous proposals. This creates possibility of an overlap in the two sets of state bits determining two nearby keystream bits. Such overlap has the potential to cancel out terms and reduce the overall “complexity”. Our choice of tap positions is motivated by the requirement of ensuring that such cancellations do not take place. The generation of a keystream bit requires a call to the majority function. We set one of the goals of choosing the tap positions to be to ensure that the inputs to the two calls to majority for generating two different keystream bits have only a small overlap. In particular, we prove that for our choice of tap positions, there is no cancellation of terms of the majority function corresponding to two keystream bits. We evolved this design criterion for the tap positions in response to an attack [6] on an earlier version, and based on informal attack ideas (though not actual attacks) suggested by Subhadeep Banik, Willi Meier, and Bin Zhang also on a previous version. Further, we choose all the tap positions for the variables in $\mathbf{X}$ to the left of all the tap positions for the variables in $\mathbf{Y}$. This choice combined with the choice of $\pi$ as the bit reversal permutation allows us to prove that the quadratic terms arising from the application of the filtering function to the state bits corresponding to two different states do not cancel out with each other. More generally we introduce a new idea, which we call the shift overlap minimisation strategy, of selecting tap positions, the principle being to try and minimise the maximum overlap that arises due to shifts.
@@ 2 Preliminaries
The fast algebraic attack (FAA) was introduced in [22]. The idea of the attack is based on the following observation. Let $f$ be an $n$-variable function and suppose $g$ is another $n$-variable function of degree $e$ such that $gf$ has degree $d$. If both $e$ and $d$ are small, then $f$ is susceptible to an FAA. Given $f$, and for $e$ and $d$ satisfying $e+d\geq n$, it is known [22] that there exists functions $g$ and $h$ with ${\sf deg}(g)=e$ and ${\sf deg}(h)\leq d$ such that $gf=h$. For any pair of functions $g$ and $h$ of degrees $e$ and $d$ respectively, satisfying $gf=h$, we have $h=gf=gf^{2}=(gf)f=hf$ and so if $h\neq 0$, then $(1\oplus h)$ is an annihilator of $f$ which implies $d\geq{\sf AI}(f)$. If $1\leq e<{\sf AI}(f)$, then $g$ is not an annihilator of $f$ and so $h=gf\neq 0$, and we obtain $e+d\geq 1+{\sf AI}(f)$. The fast algebraic immunity (FAI) of $f$ is a combined measure of resistance to both algebraic and fast algebraic attacks: ${\sf FAI}(f)\allowbreak=\allowbreak\min\left(2{\sf AI}(f),\allowbreak\min_{g\neq 0}\allowbreak\{{\sf deg}(g)\allowbreak+\allowbreak{\sf deg}(fg):\allowbreak 1\allowbreak\leq\allowbreak{\sf deg}(g)\allowbreak<\allowbreak{\sf AI}(f)\}\right)$. For any function $f$, $1+{\sf AI}(f)\leq{\sf FAI}(f)\leq 2\,{\sf AI}(f)$.
@@ 3 Construction from Maiorana-McFarland Bent Functions
Suppose that $\pi$ is an affine map, i.e. for $1\leq i\leq m$, $\pi_{i}(X_{1},\ldots,X_{m})$ is an affine function. Then
@@
Suppose now that $\pi$ is an affine map. Then $\ell(\mathbf{X})=\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle$ is an affine function. If $w=0$, then $\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})=h(\mathbf{X})$, and so we have the result. So suppose $w>0$. From Lemma 1 of [16], we have ${\sf AI}(\ell(\mathbf{X})\oplus h(\mathbf{X}))\geq{\sf AI}(h(\mathbf{X}))-1$. So
@@
From Theorem 2, choosing $\pi$ to be any affine permutation results in the AI of ${\sf MM}_{2m}$ to be lower bounded by the AI of $h$. We make the following concrete choices.
@@ 4.1 Effect of Shifts on Tap Positions
Note that both $\nu$ and $\delta$ denote the maximum overlaps (of different kinds) that can arise due to shifts. The idea of tap position selection that we put forward is to minimise $\nu$ and $\delta$. While minimising both simultaneously is difficult, we call the general principle of minimising the maximum overlap due to shifts to be the shift overlap minimisation (SOM) strategy of selecting tap positions. Selecting the tap positions to be a full positive difference set is a particular case of the SOM principle, where the minimum overlap is zero. Since using tap positions which form a full positive difference set leads to the length of the shift register being too long, the SOM principle mandates that the maximum overlap due to shifts is as small as possible.
@@ 4.2 Security Analysis
The sequence $w_{0},w_{-1},w_{-2},\ldots$ is known. If we guess the values of $s_{L-1},\ldots,s_{L-2\kappa+1}$ (a total of $\chi$ bits), then we obtain $s_{L-2\kappa}$ from the first equation, using the obtained value of $s_{L-2\kappa}$ in the second equation, we obtain the value of $s_{L-2\kappa-1}$, using the obtained value of $s_{L-2\kappa-1}$ in the third equation, we obtain the value of $s_{L-2\kappa-2}$, and so on. In other words, by guessing the values of $\chi$ of the state bits, we can obtain values of $L-\chi$ previous state bits, giving us the values of all $L$ bits of the state at a point $L-\chi$ steps in the backward direction. Once the complete state is known, the original guess can be verified by generating the keystream in the forward direction and matching with the obtained keystream. The method requires $2^{\chi}$ guesses, and for each guess about $Lm$ operations are required. By our choice of the tap positions for $\mathbf{X}$ and $\mathbf{Y}$, the value of $\chi$ is $12\kappa-$ 1. So the number of guesses required to successfully mount the attack is far greater than $2^{\kappa}$. This proves the resistance of the stream cipher to the inversion attack. Note that the inversion attack is prevented due to $\chi$ being sufficiently large. This was one of the countermeasures to the inversion attack that was already proposed in [27].
@@
Let as above the state of the LFSR at some point be $\mathbf{s}^{(0)}$ and the subsequent states of the LFSR be denoted as $\mathbf{s}^{(1)},\mathbf{s}^{(2)},\ldots,\mathbf{s}^{(c-1)}$. For $i=0,\ldots,c-1$, let $\mathbf{t}^{(i)}={\sf proj}(\mathbf{s}^{(i)})$ and $w_{i}=f_{2m+1}(\mathbf{t}^{(i)})$. Recall from the definition of $\delta$ that for $10\leq i<j\leq c-$ 1, $\mathbf{t}^{(i)}$ and $\mathbf{t}^{(j)}$ have an overlap of at most $\delta$ bits. So in the $c$ $n$-bit strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ there are at least $nc-\delta c(c-1)/2$ unknown bits and hence there are at least $2^{nc-\delta c(c-1)/2}$ possibilities for these $c$ strings. Since $f_{2m+1}$ is balanced, knowledge of $w_{0}$ reduces the number of possibilities for $\mathbf{t}^{(0)}$ by a factor of half, and more generally, the knowledge of $w_{0},\ldots,w_{c-1}$ reduces the number of possibilities for $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ by a factor of $2^{c}$. So at least $2^{(n-1)c-\delta c(c-1)/2}$ possibilities remain for the $c$ strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$. For each of these possibilities, we obtain a system of $L$ equations in $L$ variables, the solution of which provides the value of initial state of the LFSR. The correctness of the obtained value of the state can be determined by generating the keystream from the obtained value and matching with the actual keystream. Solving the system of $L$ linear equations requires $L^{3}$ arithmetic operations, and generating the keystream from the obtained state require about $Ln$ operations. So the complexity of the attack is at least $2^{(n-1)c-\delta c(c-1)/2}(L^{3}+Ln)$ operations. Suppose the following condition holds.
@@
For a filtering function $f$ having ${\sf AI}(f)=a$, an algebraic attack requires about $N=\sum_{i=0}^{a}{L\choose i}$ keystream bits and has time complexity $T$ to be about $N^{\omega}$, where $\omega$ is the exponent of matrix multiplication (see Section 3.1.5 of [14]). We assume $\omega=2.8$. For our proposal, the filtering function is $f_{2m+1}$ whose algebraic resistance is the same as that of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})$. Let $a={\sf(}f_{2m+1})={\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})$. From Proposition 4, we assume that $a=\lceil m/2\rceil$, i.e. the actual algebraic immunity of $f_{2m+1}$ is equal to the lower bound on the algebraic immunity. Let
@@
For any $(e,d)$ for which there are functions $g$ and $h$ of degrees $e$ and $d$ respectively such that $gf_{2m+1}=h$, the Berlekamp-Massey step in a fast algebraic attack on the filter generator model has time complexity $\mathcal{O}(ED\log D)$, where $E=\sum_{i=0}^{e}{L\choose i}$ and $D=\sum_{i=0}^{d}{L\choose i}$ (see [31] and Section 3.1.5 of [14]). This complexity dominates the overall time complexity of a fast algebraic attack. Ignoring the logarithm term, we take $T$ to be equal to $ED$. The number $N$ of keystream bits required is about $2E$. The maximum value of $e$ is one less than the algebraic immunity $a$. From Proposition 4, we again assume as above that $a=\lceil m/2\rceil$. Recall that for any Boolean function, its fast algebraic immunity is at least one more than its algebraic immunity, i.e. ${\sf FAI}(f_{2m+1})\geq a+1$. For any functions $g$ and $h$ of degrees $e$ and $d$ respectively such that $gf_{2m+1}=h$, we have $e+d\geq{\sf FAI}(f_{2m+1})\geq a+1$, and we assume that the lower bound is the actual value of $e+d$, and so $d=a+1-e$. Define
$\gamma$ $=$ $\min_{\stackrel{{\scriptstyle 1\leq e\leq a-1,}}{{d=a+1-e}}}\left(\sum_{i=0}^{e}{L\choose i}\right)\left(\sum_{i=0}^{d}{L\choose i}\right).$ (25)
@@ 5 Efficiency of Computing ${\sf MM}_{n}$
Using the above algorithm, it is easy to show that the computation of the weight of an $m$-bit string, when $m=2^{r}-1$ with $r\geq 1$, requires $12^{r}-r-$ 1 full adders. From this it easily follows that the number of full adders required to compute the weight of an $m$-bit string for arbitrary $m$ is $O(m)$. Since a full adder can be implemented using a constant number of NAND gates, the number of NAND gates required to compute the weight of an $m$-bit string is also $O(m)$. We record this fact in the following result.
Proposition 7.
The computation of ${\sf Maj}_{m}$ can be done using $O(m)$ NAND gates.